OpenStudy (anonymous):

I am so stuck please help 1/2 log(3)x+2log(3)y–4log(3)z??

OpenStudy (helder_edwin):

what do u have to do with that?

OpenStudy (anonymous):

find the simplified version

OpenStudy (helder_edwin):

ok. it is \[\large \frac{1}{2}\log_3x+2\log_3y-4\log_3z \] right?

OpenStudy (anonymous):

yes

OpenStudy (helder_edwin):

\[\large \frac{1}{2}\log_3x+2\log_3y-4\log_3z= \log_3x^{1/2}+\log_3y^2-\log_3z^4 \] \[\large =\log_3\left(\frac{\sqrt{x}\cdot y^2}{z^4}\right) \]

OpenStudy (anonymous):

is that the same as \[\log_{3} (√xy^2/z^4)\]

OpenStudy (helder_edwin):

yes

OpenStudy (anonymous):

thank you!

OpenStudy (helder_edwin):

u r welcome